EduFarming
Your list · 0हिं

Type at least two letters. Press Esc to close.

↑↓ navigate↵ openTAB cycle filter

Probability · Question 56 of 70

Question 56

EasyOriginal question

The probability that it rains tomorrow is 38\frac{3}{8}. What is the probability that it does not rain?

Options for question 56

Show the answer and explanation
CorrectNot quiteAnswer D

Correct answer: D

Explanation

  1. An event and its complement together are certain: P(E)+P(E′)=1P(E) + P(E') = 1.

  2. So P(E′)=1−38=58P(E') = 1 - \frac{3}{8} = \frac{5}{8}.

Concept
The complement rule
Typical time
60 s
Your time
—
The rule this testsThe addition and complement rules
Practise this chapter →

Ten questions from Probability, at your own pace.

Practise this chapter →

Ten questions from Probability, at your own pace.